# IndiTag: An Online Media Bias Analysis and Annotation System Using Fine-Grained Bias Indicators

### v1

@@ Abstract
In the age of information overload and polarized discourse, understanding media bias has become imperative for informed decision-making and fostering a balanced public discourse. This paper presents IndiTag, an innovative online media bias analysis and annotation system that leverages fine-grained bias indicators to dissect and annotate bias in digital content. IndiTag offers a novel approach by incorporating large language models, bias indicator, vector database to automatically detect and interpret bias. Complemented by a user-friendly interface facilitating both automated bias analysis and manual annotation, IndiTag offers a comprehensive platform for in-depth bias examination. We demonstrate the efficacy and versatility of IndiTag through experiments on four datasets encompassing news articles from diverse platforms. Furthermore, we discuss potential applications of IndiTag in fostering media literacy, facilitating fact-checking initiatives, and enhancing the transparency and accountability of digital media platforms. IndiTag stands as a valuable tool in the pursuit of fostering a more informed, discerning, and inclusive public discourse in the digital age. The demonstration video can be accessed from https://youtu.be/Gt2T4T7DYqs. We release an online system for end users and the source code is available at https://github.com/lylin0/IndiTag.
IndiTag: An Online Media Bias Analysis and Annotation System Using Fine-Grained Bias Indicators
[…]
@@ 1 Introduction
In an era dominated by online media, understanding and mitigating bias within digital content Yu et al. (2008); Iyyer et al. (2014); Liu et al. (2022a) have become paramount. The rapid dissemination of information through digital platforms has not only facilitated access to a wealth of knowledge but has also underscored the importance of critically evaluating the biases inherent in the content we consume. In response to these challenges, our work introduces IndiTag, a sophisticated system designed to empower users with tools for analyzing and annotating bias in online media content.
@@
In brief, by providing users with powerful tools for bias analysis and annotation, IndiTag represents a significant advancement in promoting transparency and accountability in digital media discourse. Our work underscores the importance of equipping users with the means to critically evaluate the biases present in online media content, empowering them to make informed assessments and contribute to a more nuanced understanding of digital information.
@@ 3.1 Functionality Overview
Figure 1: Functionality interpretation of IndiTag Online Analysis and Annotation System.
The IndiTag system is designed to empower users with a comprehensive suite of tools for analyzing and annotating bias in online media content. At its core, the system offers automated bias detection using advanced large language models and fine-grained bias indicators, alongside a user-friendly interface for manual annotation and validation. Users can upload digital media content and access detailed visualizations and explanations regarding bias within the content. Fig. 1 illustrates the key functionalities offered by the IndiTag system, including:
@@
  • Support for Bias Annotation and Results Downloads: In addition to automated analysis, IndiTag offers robust support for bias annotation. Users can upload documents and engage in online annotation, leveraging the system’s analysis of descriptors and indicators to effectively annotate bias within the content. Following annotation, the system compiles annotated articles, including detailed matched descriptors, indicators, the relationship between descriptors and segments, along with comments from annotators, for convenient downloading.
@@ 3.2 Workflow Overview
Figure 2: Interpretation of detailed modules of IndiTag Online Analysis and Annotation System.
As illustrated in Fig. 2, our IndiTag system comprises two stages: offline construction of a bias indicator vector database and online bias analysis and annotation. We briefly introduce these stages below to provide readers with a rough understanding of how our system operates.
@@ 4.1 Datasets
We primarily conducte experiments on political bias datasets due to their higher visibility and abundance. Our bias indicator vector database’s construction utilizes the FlipBias dataset Chen et al. (2018), sourced from allsides.com in 2018, comprising 2,781 events with articles representing different political leanings.
@@ 4.3 Evaluation and Validation
The IndiTag system undergoes thorough evaluation to ensure its effectiveness and reliability in analyzing and annotating bias in online media content. Evaluation metrics such as precision, recall, and F1-score are utilized to assess the system’s performance, providing insights into its utility for users. Our evaluation of model performance encompasses two key aspects:
@@
This evaluation framework ensures that the IndiTag system delivers reliable and accurate analysis and annotation of bias in online media content, thereby empowering users to make informed assessments and interpretations.
@@ 4.4 Experimental Results
The comparison results of bias prediction are reported in Table 1. Across all datasets, the IndiTag generally outperforms other baselines, showcasing its robustness in bias prediction tasks. Notably, IndiTag achieves competitive scores in terms of precision, recall, and F1 scores, indicating its effectiveness in identifying biased instances while maintaining a balance between precision and recall. Moreover, when considering the Micro F1 and Macro F1 scores for both biased and non-biased instances, the IndiTag demonstrates superior performance compared to other baselines. This suggests that IndiTag not only excels in detecting bias but also maintains high prediction capabilities across diverse instances, showcasing its versatility and reliability in bias prediction tasks.
@@ 5 User Interface
We present key pages of our IndiTag interface for users in Fig. 3 for reference. Fig. 3(a) demonstrates that IndiTag accommodates both single article input and structured document uploading. Fig. 3(b) and Fig. 3(c) display the generated descriptors and matched indicators for each descriptor in the input article. The bar beneath each descriptor represents the distribution of probabilities for “left”, “neutral”, and “right” leanings. Additionally, the similarity between each indicator and descriptor is visually represented to aid users in bias analysis. Fig. 3(d) introduces an additional feature enabling users to input personalized descriptors and articles to analyze the article from the user’s perspective (i.e., descriptor).
@@ 6 Conclusion
In conclusion, IndiTag represents a significant advancement in the analysis and annotation of bias in online media content. Through a meticulous two-stage approach, our system provides users with a comprehensive and accessible analysis of bias, leveraging advanced techniques from natural language processing and machine learning. By validating the effectiveness of our matched indicators through experiments across diverse datasets, we have demonstrated the meaningfulness of our approach in predicting bias labels. With its intuitive interface and powerful tools for bias analysis, IndiTag empowers users to critically evaluate digital content, contributing to transparency and accountability in digital media discourse.
@@
Our work on IndiTag has the potential to have a significant broader impact across various domains. By providing users with powerful tools for analyzing and annotating bias in online media content, our system promotes transparency and accountability in digital media discourse. This has implications for media literacy, enabling individuals to critically evaluate digital content and make informed decisions. Furthermore, IndiTag can empower journalists, researchers, and policymakers to identify and address biases within media narratives, fostering a more nuanced understanding of complex social issues. Additionally, our commitment to inclusivity and fairness ensures that IndiTag can be applied across diverse cultural and linguistic contexts, making it accessible to a global audience. Overall, our work has the potential to contribute to a more informed and equitable digital media landscape, ultimately promoting democratic values and fostering constructive dialogue in society.

### v2

# IndiTag: An Online Media Bias Analysis System Using Fine-Grained Bias Indicators for News Consumers

@@ Abstract
In the age of information overload and polarized discourse, understanding media bias has become imperative for informed decision-making and fostering a balanced public discourse. However, without the experts’ analysis, it is hard for the readers to distinguish bias from the news articles. This paper presents IndiTag, an innovative online media bias analysis system that leverages fine-grained bias indicators to dissect and distinguish bias in digital content. IndiTag offers a novel approach by incorporating large language models, bias indicators, and vector database to detect and interpret bias automatically. Complemented by a user-friendly interface facilitating automated bias analysis for readers, IndiTag offers a comprehensive platform for in-depth bias examination. We demonstrate the efficacy and versatility of IndiTag through experiments on four datasets encompassing news articles from diverse platforms. Furthermore, we discuss potential applications of IndiTag in fostering media literacy, facilitating fact-checking initiatives, and enhancing the transparency and accountability of digital media platforms. IndiTag stands as a valuable tool in the pursuit of fostering a more informed, discerning, and inclusive public discourse in the digital age. The demonstration video can be accessed from https://youtu.be/3Tux8CW46OE. We release an online system for end users and the source code is available at https://github.com/lylin0/IndiTag.
IndiTag: An Online Media Bias Analysis System Using Fine-Grained Bias Indicators for News Consumers
Luyang Lin1,2, Lingzhi Wang3thanks:   Lingzhi Wang is the corresponding author., Jinsong Guo4, Jing Li5, Kam-Fai Wong1,2 1The Chinese University of Hong Kong, China 2MoE Key Laboratory of High Confidence Software Technologies, China 3Harbin Institute of Technology, Shenzhen, China 4University College London, UK 5Department of Computing, The Hong Kong Polytechnic University, China 1,2{lylin, kfwong}@se.cuhk.edu.hk, 3wanglingzhi@hit.edu.cn 4jinsong.guo@ucl.ac.uk,5jing-amelia.li@polyu.edu.hk
@@ 1 Introduction
In an era dominated by online media, understanding and mitigating bias within digital content Yu et al. (2008); Iyyer et al. (2014); Liu et al. (2022a) have become paramount. The rapid dissemination of information through digital platforms has not only facilitated access to a wealth of knowledge but has also underscored the importance of critically evaluating the biases inherent in the content we consume. In response to these challenges, our work introduces IndiTag, a sophisticated system designed to empower news consumers with tools for analyzing bias in online media content, which only be done by journalism experts in the past.
@@
In brief, by providing news consumers with powerful tools for bias analysis, IndiTag helps to build the logistical thinking ability of readers and represents a significant advancement in promoting transparency and accountability in digital media discourse. Our work underscores the importance of equipping users with the means to critically evaluate the biases present in online media content, empowering them to make informed assessments and contribute to a more nuanced understanding of digital information.
@@ 2 Related Work
Although many experts try to organize and categorize biased articles for the readers, like AllSides 111www.allsides.com, Ad Fontes Media 222adfontesmedia.com, and Media bias / Fact Check 333mediabiasfactcheck.com, however, the analyzed articles are limited and lack of timeliness, which diminishes the role in helping the readers to distinguish the bias. To fill in this gap, IndiTag is developed for the news consumers to used directly, so that the readers can analyze the articles they are reading timely.
@@ 3.1 Functionality Overview
Figure 1: Functionality interpretation of IndiTag Online Analysis and Note-taking System.
The IndiTag system is designed to empower users with a comprehensive suite of tools for analyzing and distinguishing bias in online media content. At its core, the system offers automated bias detection using advanced large language models and fine-grained bias indicators, alongside a user-friendly interface for news consumers to record and validate the results. Users can upload digital media content and access detailed visualizations and explanations regarding bias within the content. Fig.˜1 illustrates the key functionalities offered by the IndiTag system, including:
@@
  • Support for Notes-Taking and Results Downloads: In addition to automated analysis, IndiTag offers support for the news consumers to take reading notes. Users can upload documents and engage in online bias note-taking, leveraging the system’s analysis of descriptors and indicators to effectively distinguish bias within the content. Following adding the notes, the system compiles the analyzed articles, including detailed matched descriptors, indicators, and the relationship between descriptors and segments, along with comments from users, for convenient downloading.
@@ 3.2 Workflow Overview
Figure 2: Interpretation of detailed modules of IndiTag Online Analysis and Note-taking System.
As illustrated in Fig.˜2, our IndiTag system comprises two stages: offline construction of a bias indicator vector database and online bias analysis and note-taking. We briefly introduce these stages below to provide readers with a rough understanding of how our system operates.
@@ 4.1 Datasets
We primarily conduct the system evaluation on political bias datasets due to their higher visibility and abundance. Our bias indicator vector database’s construction utilizes the FlipBias dataset Chen et al. (2018), sourced from allsides.com in 2018, comprising 2,781 events with articles representing different political leanings.
@@ 4.3 Evaluation and Validation
The IndiTag system undergoes thorough evaluation to ensure its effectiveness and reliability in analyzing bias in online media content. Evaluation metrics such as precision, recall, and F1-score are utilized to assess the system’s performance, providing insights into its utility for users. Our evaluation of model performance encompasses two key aspects:
@@
This evaluation framework ensures that the IndiTag system delivers reliable and accurate analysis bias aligned with humans in online media content, thereby empowering users to make informed assessments and interpretations.
@@ 4.4 Experimental Results
The comparison results of bias prediction are reported in Table˜1. Across all datasets, the IndiTag generally outperforms other baselines, showcasing its robustness in bias prediction tasks. Notably, IndiTag achieves competitive scores in terms of precision, recall, and F1 scores, indicating its effectiveness in identifying biased instances while maintaining a balance between precision and recall. Moreover, when considering the Micro F1 and Macro F1 scores for both biased and non-biased instances, the IndiTag demonstrates superior performance compared to other baselines. This suggests that IndiTag not only excels in detecting bias but also maintains high prediction capabilities across diverse instances, showcasing its versatility and reliability in bias prediction tasks.
@@ 5 User Interface
We present key pages of our IndiTag interface for users in Fig.˜3 for reference. Fig.˜3(a) demonstrates that IndiTag accommodates both single article input and structured document uploading. Fig.˜3(b) and Fig.˜3(c) display the generated descriptors and matched indicators for each descriptor in the input article. The bar beneath each descriptor represents the distribution of probabilities for “left”, “neutral”, and “right” leanings. Additionally, the similarity between each indicator and descriptor is visually represented to aid users in bias analysis. Fig.˜3(d) introduces an additional feature enabling users to input personalized descriptors and articles to analyze the article from the user’s perspective (i.e., descriptor).
@@ 6 Conclusion
In conclusion, IndiTag represents a significant advancement in analyzing the bias of online media content for news consumers. Through a meticulous two-stage approach, our system provides users with a comprehensive and accessible analysis of bias, leveraging advanced techniques from natural language processing and machine learning. By validating the effectiveness of our matched indicators through evaluating across diverse datasets, we have demonstrated the meaningfulness of our approach in predicting bias labels. With its intuitive interface and powerful tools for bias analysis, IndiTag empowers users to critically evaluate digital content, contributing to transparency and accountability in digital media discourse.
@@
Our work on IndiTag has the potential to have a significant broader impact across various domains. By providing users with powerful tools for analyzing bias in online media content, our system promotes transparency and accountability in digital media discourse. This has implications for media literacy, enabling individuals to critically evaluate digital content and make informed decisions. Furthermore, IndiTag can empower journalists, researchers, and policymakers to identify and address biases within media narratives, fostering a more nuanced understanding of complex social issues. Additionally, our commitment to inclusivity and fairness ensures that IndiTag can be applied across diverse cultural and linguistic contexts, making it accessible to a global audience. Overall, our work has the potential to contribute to a more informed and equitable digital media landscape, ultimately promoting democratic values and fostering constructive dialogue in society.